OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 5x + y = 1 3x + 2y = 2

OpenStudy (immanuelv):

5x + y = 1 3x + 2y = 2 for that you need to isolate y in first equation y = -5x +1 then plug in that y to the second equation 3x + 2y = 2 3x + 2(-5x+1) = 2 3x+ -10x+2 = 4 3x - 10x +2 =4 -7x = 2 x = -2/7

OpenStudy (immanuelv):

then plug in -2/7 for x to get y 5x + y =1 5(-2/7) +y = 1 -10/7 +y =1 y = 17/7

OpenStudy (anonymous):

Thank you

OpenStudy (immanuelv):

you're welcome :)